OpenStudy (anonymous):

Put in point slope form and slope intercept form y1=14% y2=12.5% x1=5 x2=13

OpenStudy (gorv):

we are give 2 points here

OpenStudy (gorv):

14%=14/100 =0.14

OpenStudy (gorv):

12.5%=12.5/100=0.125

OpenStudy (gorv):

(x1,y1) and (x2,y2)

OpenStudy (gorv):

can u do it now

OpenStudy (gorv):

you have 2 points

OpenStudy (gorv):

point slope form y-y1=((y2-y1)/(x2-x1))*(x-x1)

OpenStudy (gorv):

slope =(y2-y1)/(x2-x1)

OpenStudy (gorv):

slope intercept form y=mx+c

OpenStudy (gorv):

m=slope

OpenStudy (gorv):

try it now

OpenStudy (anonymous):

whoops sorry went and got my laundry. umm i dont know. I got y-14 for the first side of point slope form and that was right

OpenStudy (anonymous):

oh wait i see hold on

OpenStudy (anonymous):

okay i dont know